# AgentA/B: Automated and Scalable Web A/B Testing with Interactive LLM Agents

### v2

@@ 5.3. Findings: System Effectiveness Across Interface Variants
To assess whether AgentA/B can detect subtle differences between design variants, we compared LLM agent behavior across control and treatment conditions featuring different filter configurations. As shown in Table 5.1, agents exhibited notable condition-dependent behavioral changes in response to the design manipulation.
@@
Agents in the treatment condition clicked on more products (2.09 vs. 1.87), and the sessions have a higher total number of actions (6.60 vs. 6.05, t(998)=1.08, p=0.28). Most notably, the LLM agents in the treatment condition click filter options more frequently than the ones in the control condition (0.60 vs. 0.58, t(998)=1.00, p=0.32), which suggests that the treatment design that removes unrelated filters improved filter discoverability. Lastly, when looking at the final outcome of the sessions in the treatment condition and control condition, LLM agent simulated virtual customers purchased more products (414 v.s. 404, $\chi^{2}(1)=5.51$, p=0.03) , and spent more money (\$60.99 v.s. \$55.14, t(998)=0.39, p=0.69) than the virtual customers in the control condition. Most of these comparisons, except the “# of product purchase” are not statistically significant, which suggests to the A/B test owner that this filter interface design demands a higher number of participants size (N). Importantly, the LLM-agent A/B testing results are consistent with the direction of changes in the human A/B testing results (which can not reported due to privacy policy).
@@
In summary, the A/B simulation conducted with AgentA/B reveals that LLM-based agents, when equipped with structured personas and embedded intentions, produce behaviorally rich and interface-sensitive interactions. These behaviors vary in predictable ways based on persona, align with human user baselines, and are capable of distinguishing between subtle interface design variants. These findings validate the system’s use as a behaviorally grounded, scalable alternative to conventional A/B testing, suitable for early-stage prototyping, pre-deployment design validation, and hypothesis-driven UX evaluation.
@@ 6. Discussion
6.1. Accelerating Design Iteration Through Simulated Feedback Loops
@@ Appendix A Example of Persona Generated by AgentA/B
{minted}
[breaklines, breaksymbol=,breakanywhere]text Persona: Marcus
Background: Marcus is a 35-year-old freelance graphic designer living in Austin, Texas. After working for a decade in various creative agencies, he transitioned to freelancing to gain more control over his schedule and focus on passion projects, such as illustrating indie game assets and creating digital art for local musicians.
Demographics:
Age: 35
Gender: Male
Education: Bachelor’s degree in Visual Communication
Profession: Freelance Graphic Designer
Income: $70,000(variablebasedonprojects)\par FinancialSituation:Marcusearnsadecentlivingfromhisfreelancegigs,thoughhisincomecanfluctuate.He^{\prime}sfinanciallystablebutcautiousaboutbigexpenses.Hesetsasidepartofhisearningsfortravelandsoftwareupgrades,whichareessentialforhiswork.\par ShoppingHabits:Marcusenjoysdiscoveringuniqueornicheproducts,especiallytechgadgets,artsupplies,andstreetwear.Heprefersshoppingonlineforthevarietyandreadsreviewscarefully.He^{\prime}sbrand-loyalwhenitcomestotoolsherelieson,likehisdrawingtabletanddesignsoftware.Forclothing,he^{\prime}sdrawntobold,graphic-heavyitemsthatreflecthisartisticvibe.\par ProfessionalLife:Marcusworksfromahomestudiothatdoublesasacreativespace.Hecollaboratesremotelywithclientsfromdifferentindustries,jugglesmultipledeadlines,andoftenpullslatenights.Hefrequentlyupdateshisportfolioandmaintainsastrongsocialmediapresencetoattractnewclients.\par PersonalStyle:Marcushasanedgyandexpressivefashionsense.Hewearslarge-sizedclothingandgravitatestowarddarktoneswithpopsofneonorgraphicprints.Comfortisimportant,buthelikestomakeavisualstatementwithwhathewears.Hisgo-tooutfitisasofthoodiewithacustomdesign,blackjoggers,andhigh-topsneakers.\par\@add@PDF@RDFa@triples\par$

### v3

@@ 5.3. Findings: System Effectiveness Across Interface Variants
To assess whether Agent A/B can detect subtle differences between design variants, we compared LLM agent behavior across control and treatment conditions featuring different filter configurations. As shown in Table 2, agents exhibited notable condition-dependent behavioral changes in response to the design manipulation.
@@
Agents in the treatment condition clicked on more products (2.09 vs. 1.87), and the sessions have a higher total number of actions (6.60 vs. 6.05, t(998)=1.08, p=0.28). Most notably, the LLM agents in the treatment condition click filter options more frequently than the ones in the control condition (0.60 vs. 0.58, t(998)=1.00, p=0.32), which suggests that the treatment design that removes unrelated filters improved filter discoverability.
When looking at the final outcomes of the sessions in the treatment and control conditions, LLM-agent simulated virtual customers purchased more products in the treatment condition compared to the control condition(414 v.s. 404, $\chi^{2}(1)=5.51$, p=0.03). This represents a modest but statistically significant increase in purchase completion, suggesting that reducing the number of filter options improved the likelihood of customers making a purchase. Expressed in practical terms, the treatment condition resulted in roughly a 2–3% higher conversion rate, which could translate into millions of additional purchases per day on Amazon.
In terms of average spending across all customers, LLM agent simulated virtual customers in the treatment condition spent somewhat more than those in the control condition (\$60.99 v.s. \$55.14). However, this difference was not statistically significant (t(998) = 0.39, p = 0.69). The mean difference of about \$6 is small relative to the variability in spending, suggesting that while the treatment may have had a slight positive effect on purchase amounts, substantially larger sample sizes would be required to reliably detect such an effect.
Breaking down average spending by gender, the data show different patterns across groups. Female customers spent an average of \$49.18 in the control condition and \$51.30 in the treatment condition, indicating only a small increase under the treatment. Male customers, by contrast, showed a larger increase, from \$61.82 (control) to \$70.83 (treatment). While statistical tests were not conducted for each subgroup separately, the direction of change suggests that the simplified filter interface may have a stronger positive effect on male customers’ spending than on female customers.
A clearer set of differences emerges when comparing average spending across age groups. Among younger customers ($\leq 35$ ), average spending actually decreased in the treatment condition compared to the control condition ($\mathdollar 56.68\rightarrow\mathdollar 43.41$). Middle-aged customers (35–55) spent more under treatment ($\mathdollar 59.21\rightarrow\mathdollar 71.19$), while older customers (¿55) showed the largest increase ($\mathdollar 40.24\rightarrow\mathdollar 79.07$). These results suggest that the reduced filter interface may be particularly beneficial for older customers, potentially because it helps reduce choice overload. Conversely, younger customers may rely more on detailed filtering to find products that match their preferences, leading to lower spending when filter options are reduced.
In sum, reducing the number of filter options was associated with a higher purchase completion rate and possible increases in customer spending, though the latter effects did not reach statistical significance. Subgroup analyses suggest heterogeneity in response: males and older customers appear to benefit more from the simplified interface, while younger customers may prefer more filtering flexibility. Given the variability in spending outcomes, further testing with larger samples will be important to confirm these trends.
Notably, the LLM-agent A/B testing results aligned directionally with patterns observed in human A/B testing . This convergence increases confidence that agent-based simulations can anticipate real user behaviors and provide early insight into interface performance.
@@
Overall, the A/B simulation conducted with Agent A/B shows that LLM-based agents, when guided by personas and embedded intentions, generate behaviorally rich and interface-sensitive interactions. These behaviors vary in predictable ways across personas, align with humans, and detect subtle differences in design. Such findings support the system’s role as a behaviorally grounded, scalable complement to conventional A/B testing. The results suggest Agent A/B is well-suited for early-stage prototyping, pre-deployment validation, and hypothesis-driven UX research.
@@ 6. Discussion
6.1. Accelerating Design Iteration with Simulated Feedback
@@ Appendix A Example of Persona Generated by Agent A/B
Persona: Marcus
[…]
Marcus is a 35-year-old freelance graphic designer living in Austin, Texas. After working for a decade in various creative agencies, he transitioned to freelancing to gain more control over his schedule and focus on passion projects, such as illustrating indie game assets and creating digital art for local musicians.
[…]
Marcus earns a decent living from his freelance gigs, though his income can fluctuate. Hes financially stable but cautious about big expenses. He sets aside part of his earnings for travel and software upgrades, which are essential for his work.
[…]
Marcus enjoys discovering unique or niche products, especially tech gadgets, art supplies, and streetwear. He prefers shopping online for the variety and reads reviews carefully. Hes brand-loyal when it comes to tools he relies on, like his drawing tablet and design software. For clothing, hes drawn to bold, graphic-heavy items that reflect his artistic vibe.
[…]
Marcus works from a home studio that doubles as a creative space. He collaborates remotely with clients from different industries, juggles multiple deadlines, and often pulls late nights. He frequently updates his portfolio and maintains a strong social media presence to attract new clients.
[…]
Marcus has an edgy and expressive fashion sense. He wears large-sized clothing and gravitates toward dark tones with pops of neon or graphic prints. Comfort is important, but he likes to make a visual statement with what he wears. His go-to outfit is a soft hoodie with a custom design, black joggers, and high-top sneakers.